OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, -5) with a slope of -1?

OpenStudy (tyteen4a03):

Plug the numbers in the point-slope form: \(y-y1 = m(x-x1)\).

OpenStudy (anonymous):

thanks tyteen i knew it was something like that :D

OpenStudy (anonymous):

Is it y = -x-4 ?

OpenStudy (anonymous):

no it is x+y=-4

OpenStudy (anonymous):

is that "standard form" ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

what "form" is y = -x-4 ?

OpenStudy (anonymous):

its in slope-intercept form

OpenStudy (anonymous):

gotcha, thanks for explaining :)

OpenStudy (anonymous):

ur welcome